# DERD-Net: Learning Depth from Event-based Ray Densities

### v1

@@ 1 Introduction
Unlike conventional cameras, event cameras operate asynchronously, detecting per-pixel brightness changes [22, 8, 9]. This provides high temporal resolution and robustness to motion blur, making them particularly suited for dynamic scenes. Their sparse output enables efficient processing of only relevant areas, making them ideal for real-time tasks such as visual odometry (VO) / SLAM.
@@
Harnessing deep learning for depth estimation with event cameras has the potential to transform such applications. However, adapting neural networks to event data remains challenging due to its asynchronous stream-like nature. Moreover, the scarcity of event camera datasets with ground truth depth [48, 14] results in limited training data, which can lead to overfitting [38]. While simulating data is one way to address this issue, generalizing models trained on simulation to real-world scenarios is far from trivial due to differences in data distribution [27, 36]. Thus, instead of directly processing events (which is prone to overfitting and extensive training time) we rely on intermediate event representations informed by the scene geometry.
@@
One promising approach to capture the 3D scene structure is back-projecting events as rays into space and capturing their intersection densities as disparity space images (DSIs) [31]. DSIs from two or more cameras can be fused, eliminating the need for event synchronization between cameras. This reduces complexity and allows for more robust depth estimation. The Multi-Camera Event-based Multi-View Stereo (MC-EMVS) method has recently produced state-of-the-art (SOTA) results, outperforming other techniques in depth benchmarks across several metrics [13]. To obtain pixel-wise depth estimations from a DSI, MC-EMVS selects the disparity level with the highest ray count, effectively using an argmax operation. A selective threshold filter is applied to predict depth only for pixels with sufficient ray counts.
@@
While straightforward, this approach does not fully utilize the potential of DSIs. It is susceptible to noise and limited in the information it can extract due to ignoring the surrounding pixels and more complex patterns across disparity levels. Consequently, it leads to fewer pixels obtaining depth estimation and less accurate depth predictions than the DSI might potentially allow for. Therefore, there is a need for a more effective approach that can increase the reliability, depth accuracy, and number of pixels for which depth is estimated by recognizing and adequately processing more complex data patterns within the DSIs.
@@
Our Contribution. We propose a novel deep learning framework for event-based depth estimation that is optimized for SLAM scenarios and addresses the aforementioned limitations. An overview is illustrated in Fig. 1. Our approach estimates pixel-wise depth from a DSI using a neural network with 3D convolutions and a recurrent structure. The framework is directly applicable to both monocular and stereo settings. Our key contributions include:
@@
Learning-based Local Processing: For each selected pixel, a small local subregion of the DSI (Sub-DSI) is used as input to the neural network. This novel design leverages the inherent sparsity of event data and DSIs to efficiently process and produce only relevant information for sparse depth-related tasks. (Sec. 3).
@@
Enhanced Data Utilization: Our model captures complex patterns within the DSI, increasing depth prediction accuracy and the number of pixels for which depth can be reliably estimated. Limiting the input to small local subregions around selected pixels enhances generalization by preventing the network from overfitting to specific scenes and augmenting the available training set, as each Sub-DSI serves as an individual data instance.
@@
Efficiency and Scalability: By adopting our Sub-DSI approach, we obtain small independent inputs of fixed size. This enables full parallelization and provides an ultra-light network that has the ability to handle any camera resolution with constant very short inference time. Our network architecture allows the processing of DSIs of variable depth resolution. (Sec. 3.2).
@@
Comprehensive Experiments: We evaluate our model using both monocular and stereo data from standard datasets MVSEC [48] and DSEC [10], employing cross-validation, outperforming the state of the art by a large margin on ten figures of merit. When applied to monocular data, our model achieves performance comparable to SOTA methods that require stereo data (Sec. 4).
@@
To the best of our knowledge, our work is the first learning-based multi-view stereo method to (i) use events and camera poses as input, a crucial requirement for accurate depth prediction over long intervals (as used in SLAM [15]); (ii) demonstrate successful and generalizable depth prediction on real-world event data from DSIs; (iii) report good generalization on all three MVSEC indoor flying sequences [14], even including hybrid methods which combine stereo intensity frames with events.
@@ 2 Related Work
Stereo depth estimation with event cameras has been an intriguing problem since the invention of the first event camera by Mahowald and Mead in the 1990s [26, 14, 9] due to their potential for high temporal resolution and robustness to motion blur. Recent approaches have addressed stereo event-based 3D reconstruction for VO and SLAM [45, 47, 46, 7, 15, 35]. These methods assume a static world and known camera motion, using this information to assimilate events over longer time intervals, thereby increasing parallax and producing more accurate semi-dense depth maps. A comprehensive review is provided in [14].
MC-EMVS [13] introduced a novel stereo approach for depth estimation which does not require explicit data association, using DSIs generated from stereo events cameras. By leveraging the sparsity of events and fusing back-projected rays, they outperformed the event-matching based solution of [46] and thereby achieved SOTA results in 3D reconstruction and VO [15]. We advance this approach by employing a compact neural network specialized to derive explicit depth from the DSIs, creating a standardized and effective framework for processing event-based data in deep learning applications that does not rely on event simultaneity or matching.
Deep Learning for depth estimation from event data. Deep learning has significantly advanced depth estimation in traditional monocular and stereo camera setups, achieving remarkable results [21, 20, 37]. However, its application to event camera data remains relatively limited due to the continuous nature of event streams, which require specialized frameworks [14]. For example, in monocular vision, [16] uses synthetic data on a recurrent network to capture temporal information from grid-like event inputs. Yet, mismatches between synthetic and real data degrade performance [32], and monocular depth estimation from events is an ill-posed problem, making high accuracy challenging to achieve with this learning framework [44].
For stereo depth estimation, [38, 40] present two pioneering studies. Specifically, DDES [38] introduced the first deep-learning–based supervised stereo-matching method, while [40] proposed the first unsupervised learning framework. Both methods utilize a First-In First-Out queue to store events at each position, allowing for concurrent time and polarity reservation. Nonetheless, limitations for using visual cues from both left and right event cameras are imposed by the accompanying increase in model complexity and memory costs. Our method overcomes these challenges by maintaining constant input dimensions, defined by the size of the DSI subregion around a selected pixel. It thereby provides a significant advancement in applying deep learning to event-based depth estimation on real-world data.
@@ 3 Methodology
In this section, we present our supervised-learning–based approach and the related framework in detail, for which a general overview is provided in Fig. 1. We describe the preprocessing of the data, the architecture of our network (shown in Fig. 2) and the training and inference procedures.
@@ 3.1 Framework
As an event camera with pixels moves through a scene, it triggers events and produces a near time-continuous stream of data . Following [31, 13], this stream is sliced by dividing time into intervals. For every time interval, a representative camera viewpoint is selected and a DSI is created. Events are then back-projected into 3D space by casting rays from this viewpoint through the corresponding pixel . The depth axis is discretized into levels, equidistant in inverse linear space, resulting in a 3D DSI of size voxels, whose values represent the number of rays passing through each region (i.e., voxel) of space.
@@
Since DSIs are typically large and sparse, depth is estimated only for pixels with sufficient information. A confidence map is generated by projecting the DSI onto a 2D grid of size , where each pixel’s value represents the maximum ray density among all depth levels [13] (pixel selection map in Fig. 3). An adaptive Gaussian threshold (AGT) filter is then applied to this grid to select the pixels with a sufficient maximum ray density for reliable depth estimation. For each selected pixel , a surrounding subregion is extracted from the DSI, including the ray counts of all pixels within an L1 radii of :
@@
and serves as input to our neural network. Using only small subregions of the DSIs as inputs promotes an efficient and compact architecture operating independent of camera resolution and reduces the risk for overfitting.
@@ 3.2 Network Architecture
The architecture of the neural network is illustrated in Fig. 2, with the dimensions of each layer listed in Tab. 1. The network receives the normalized Sub-DSI (1) as input. The objective is to capture local geometrical patterns in the Sub-DSI to extract more relevant depth information from it than the SOTA argmax approach used in [31, 13]. Since established networks like U-Net [33] often include strong spatial compression and Transformers tend to impose high data demands for reliable generalization [24], we tailor an ultra-lightweight custom architecture that shares design elements with FireNet [34], adapted for very small frame sizes yet variable depth dimensions. Similarly to RAFT-Stereo [23], we adopt convolutions and a Gated Recurrent Unit (GRU) to efficiently handle different depth resolutions, enabling customization of the desired depth precision without modifying architecture.
@@
Finally, a dense layer that preserves the dimension of with ReLU activation and a subsequent output dense layer are applied to process the hidden state. We introduce two versions of the network for custom modification of depth estimation density. In the single-pixel version, the network predicts the depth for the selected pixel while in the multi-pixel version, the output is a grid representing the depth predictions of the central pixel and its 8 neighbors.
Layer Dimensions Details Sub-DSI (Input) Depth Channels Width Height 3D-Convolution K = (3,3,3); P = (1,0,0); S = (2,1,1) ReLU + Flattens Flatten channels and frame GRU Selects final hidden state Dense + ReLU Maintains dimension Dense (Output) or Outputs depth value(s)
@@ 3.3 Training and Inference
As supervised loss function for training the neural network model we use the mean absolute error (MAE), with given ground truth depth (this is the case of standard real-world datasets used, such as MVSEC and DSEC – see Sec. 4). To reduce training time and further improve generalization, we additionally employ ensemble learning (EL) [28, 29].
For training, we initialize two identical but independent instances of our neural network with different random weights. The training set is split into two disjoint subsets, enabling parallel training. During testing or inference, each Sub-DSI is processed simultaneously by both networks, and the final depth estimation is obtained by averaging the individual predictions. This helps reduce variance in the predictions, leading to more stable and accurate results. While we use two models, this approach is easily scalable, and incorporating additional models could further improve robustness.
Dataset Dataset Details DSI Parameters Gauss. Filter Training Process Sequences Resolution LiDAR Interval Span Min Dist Max Dist D Sub-DSI Window C Batch Optimizer LR LF Epochs MVSEC Indoor flying 1,2,3 346 260 px 50 ms 1 s 1 m 6.5 m 100 77 99 10 64 AdamW MAE 3 DSEC Zurich04a 640 480 px 100 ms 0.2 s 4 m 50 m 100 77 99 2 64 AdamW MAE 3
@@ 4 Experiments
Supervised training and testing are performed on the network. We investigate the impact of DSIs derived from both monocular and stereo settings during training and testing. Table 2 provides an overview of the key parameters used in the datasets and the training processes of the experiments.
@@ 4.1 Metrics
The performance of the networks is evaluated using ten standard metrics commonly employed in depth estimation tasks [13]. We calculate both mean and median errors between the estimated and GT depths, with median errors providing robustness against outliers. Additionally, we report the number of reconstructed points, reflecting the algorithm’s ability to generate valid depth estimations, and the number of outliers (bad-pix [11]), representing the proportion of significant depth estimation errors. In the supplementary, we also compute the scale-invariant logarithmic error (SILog Err) to evaluate the error while considering scale, and the sum of absolute relative differences (AErrR) to assess the relative accuracy of the depth predictions. Finally, we report -accuracy values, which indicate the percentage of points whose estimated depth falls within specified limits relative to GT [41].
@@ 4.2 Baseline Methods
We compare our proposed approach against several SOTA methods that have been benchmarked on the task of long-term event-based depth estimation [14], which have been evaluated under the same input conditions (events and camera poses) and output format (semi-dense depth maps) supportive of SLAM.
The Generalized Time-Based Stereovision (GTS) method [18] utilizes a two-step process: first performing stereo matching based on a time-consistency score for each event, followed by depth estimation through triangulation. The Semi-Global Matching (SGM) method [17] is adapted for event-based data by generating time images and subsequently applying stereo matching, with the depth map being refined by masking it at the locations of recent events. Another method, Event-based Stereo Visual Odometry (ESVO) [46], integrates depth estimates by employing Student-t filters, ensuring robust spatio-temporal consistency between stereo time image patches.
The two closest baseline methods for performance comparison of our method are EMVS for monocular vision [31] and MC-EMVS for stereo vision [13]. Both methods extract pixel-wise depth from DSIs by applying the argmax function. To ensure consistency and fairness, we benchmark the methods following the procedure used in prior works [14].

### v2

@@ 1 Introduction
Unlike conventional cameras, event cameras operate asynchronously, detecting per-pixel brightness changes (called “events”) [1, 2, 3]. This provides high temporal resolution and robustness to motion blur, making them well suited for dynamic scenes. Their sparse output enables efficient processing of only relevant areas, making them ideal for real-time tasks such as visual odometry (VO) / SLAM.
@@
Harnessing deep learning for depth estimation with event cameras has the potential to transform such applications. However, adapting neural networks to event data remains challenging due to its asynchronous stream-like nature. Moreover, the scarcity of event camera datasets with ground truth depth [4, 5] results in limited training data, which can lead to overfitting [6]. While simulating data is one way to address this issue, generalizing models trained on simulation to real-world scenarios is far from trivial due to differences in data distribution [7, 8, 9, 10]. Thus, instead of directly processing events (which is prone to overfitting and extensive training time) we rely on intermediate event representations informed by the scene geometry.
@@
One promising approach for 3D reconstruction is back-projecting events as rays into space and capturing their intersection densities as disparity space images (DSIs) [11]. DSIs from two or more cameras can be fused, eliminating the need for event synchronization between cameras. This reduces complexity and allows for more robust depth estimation. The Multi-Camera Event-based Multi-View Stereo (MC-EMVS) method [12] recently produced state-of-the-art (SOTA) results, outperforming other techniques in depth benchmarks across several metrics. To obtain pixel-wise depth estimates from a DSI, MC-EMVS selects the disparity level with the highest ray count, effectively using an argmax operation. Ray counting is used as a proxy for finding 3D edges where the rays intersect. A selective threshold filter is applied to predict depth only for pixels with sufficient ray counts.
@@
While straightforward, this approach does not fully utilize the potential of DSIs. It is susceptible to noise and cannot effectively extract cues from surrounding pixels and more complex patterns across disparity levels. Consequently, it leads to fewer pixels obtaining depth estimation and less accurate depth predictions than the DSI might potentially allow for. Therefore, we need more effective approaches for extracting depth from DSIs, that are more reliable, accurate, and produce more depth estimates, by adequately recognizing complex ray intersection patterns.
@@
Our Contribution. We propose a novel deep learning framework for event-based depth estimation that is optimized for SLAM scenarios and addresses the aforementioned limitations. An overview is illustrated in Fig.˜1. Our approach estimates pixel-wise depth from a DSI using a neural network with 3D convolutions and a recurrent structure. The framework is directly applicable to both monocular and stereo settings. Our key contributions include:
@@
Learning-based Local Processing: For each selected pixel, a small local subregion of the DSI (Sub-DSI) is used as input to the neural network. This novel design leverages the inherent sparsity of event data and DSIs to efficiently process and produce only relevant information for sparse depth-related tasks. (Sec.˜3).
@@
Enhanced Data Utilization: Our model captures complex patterns within the DSI, increasing depth prediction accuracy and the number of pixels for which depth can be reliably estimated. Limiting the input to small local subregions around selected pixels enhances generalization by preventing the network from overfitting to dynamics specific to the training scenes and augmenting the available training set, as each Sub-DSI serves as an individual data instance.
@@
Efficiency and Scalability: By adopting our Sub-DSI approach, we obtain small independent inputs of fixed size. This enables full parallelization and provides an ultra-light network that has the ability to handle any camera resolution with constant very short inference time. Our network architecture allows the processing of DSIs of variable depth resolution. (Sec.˜3.2).
@@
Comprehensive Experiments: We evaluate our model on both monocular and stereo data from the standard datasets MVSEC [4] and DSEC [13] using cross-validation. It outperforms the state of the art by a large margin on ten figures of merit. Even using monocular data, our model achieves performance comparable to SOTA methods that require stereo data (Sec.˜4). We show downstream applicability and robustness to imperfect (noisy) camera poses.
@@
To the best of our knowledge, our work is the first learning-based multi-view stereo method to (i) use camera poses along with events as input, which is crucial for accurate depth estimation over long intervals (as used in SLAM [14, 15]); (ii) demonstrate successful and robust depth prediction on real-world event data from DSIs; (iii) report good generalization on all three MVSEC indoor flying sequences [5], even when compared to multi-modal methods that combine stereo intensity frames with events. We provide code, trained models and video results for clarity and reproducibility.
@@ 2 Related Work
Stereo depth estimation with event cameras has been a captivating problem since the invention of the first event camera by Mahowald and Mead in the 1990s [16, 5, 3] due to their potential for high temporal resolution and robustness to motion blur. Recent approaches have addressed stereo event-based 3D reconstruction for VO and SLAM [17, 18, 19, 20, 14, 21, 22]. These methods assume a static world and known camera motion, using this information to assimilate events over longer time intervals, thereby increasing parallax and producing more accurate semi-dense depth maps. A comprehensive review is provided in [5].
MC-EMVS [12] introduced a novel stereo approach for depth estimation which does not require explicit data association, using DSIs generated from stereo events cameras. By leveraging the sparsity of events and fusing back-projected rays, they outperformed the event-matching–based solution of [19] and thereby achieved SOTA results in 3D reconstruction and VO [14]. Evidently, this DSI-based 3D reconstruction is robust to imperfect poses estimated using an event camera tracking method. We advance this approach by employing a compact neural network specialized to derive explicit depth from the DSIs, creating a standardized and effective framework for processing event-based data in deep learning applications that does not rely on event simultaneity or matching.
Deep Learning for depth estimation from event data. Deep learning has significantly advanced depth estimation in traditional monocular and stereo camera setups, achieving remarkable results [23, 24, 25]. However, its application to event camera data remains relatively limited due to the sparse asynchronous nature of event streams, which require specialized frameworks [5]. For example, in monocular vision, [26] uses synthetic data on a recurrent network to capture temporal information from grid-like event inputs. Yet, mismatches between synthetic and real data degrade performance [27], and monocular depth estimation from events is an ill-posed problem, making high accuracy challenging to achieve with this learning-based framework [28].
For stereo depth estimation, [6, 29] present two pioneering studies. Specifically, DDES [6] introduced the first deep-learning–based supervised stereo-matching method, while [29] proposed the first unsupervised learning framework. Both methods use First-In First-Out queues to store events at each position, allowing for concurrent time and polarity reservation. Nevertheless, high event rates lead to greater processing demands and, consequently, increased model complexity and memory requirements, limiting the use of visual cues from both event cameras. Our method overcomes these challenges by maintaining constant input dimensions defined by the size of the DSI subregion around a selected pixel, regardless of the event rate. It thereby provides a significant advancement in applying deep learning to event-based depth estimation on real-world data.
@@ 3 Methodology
In this section, we present our supervised-learning–based approach and the related framework in detail, for which a general overview is provided in Fig.˜1. We describe the preprocessing of the data, the architecture of our network (shown in Fig.˜2) and the training and inference procedures.
@@ 3.1 Framework
As an event camera with pixels moves through a scene, it triggers events and produces a near continuous-time stream of data . Following [11, 12], this stream is sliced into time intervals. For every time interval, a DSI is created and associated with a camera viewpoint (called Reference Viewpoint) as follows: given camera poses, all events in the interval are back-projected into 3D space by casting rays from the moving camera optical center through the corresponding pixel . The depth axis is discretized into levels, equidistant in inverse linear space, resulting in a 3D DSI of size voxels, whose values represent the number of rays passing through each region (i.e., voxel) of space (as shown in Fig.˜1). Although input poses are required for building a DSI, they can be obtained from tracking methods or dead-reckoning [30, 12, 14].
@@
Since DSIs are typically large and sparse, depth is estimated only for pixels with sufficient information. A confidence map is generated by projecting the DSI onto a 2D grid of size , where each pixel’s value represents the maximum ray density among all depth levels [12] (called pixel selection map in Fig.˜3). An adaptive Gaussian threshold (AGT) filter is then applied to this grid to select the pixels with a sufficient maximum ray density for reliable depth estimation. For each selected pixel , a surrounding subregion is extracted from the DSI, including the ray counts of all pixels within L1 radii of :
@@
and serves as input to our neural network. Using only small subregions of the DSIs as inputs leads to an efficient and compact architecture operating independently of camera resolution. Furthermore, it reduces the risk of overfitting by encouraging the model to learn generalizable patterns of ray intersections within the Sub-DSIs instead of memorizing semantics and dynamics specific to the training scene. Such localized input processing is possible because DSIs consolidate sparse events into a structured format that preserves geometric information within small spatial regions.
@@ 3.2 Network Architecture
Layer Dimensions Details Sub-DSI (Input) Depth Channels Width Height 3D-Convolution K = (3,3,3); P = (1,0,0); S = (2,1,1) ReLU + Flatten Flattens channels and frame GRU Selects final hidden state Dense + ReLU Maintains dimension Dense (Output) or Outputs depth value(s)
The architecture of the neural network is illustrated in Fig.˜2, with the dimensions of each layer listed in Tab.˜1. The network receives the normalized Sub-DSI (1) as input. The objective is to capture local geometrical patterns in the Sub-DSI to extract more relevant depth information than the SOTA argmax approach used in [11, 12]. Since established networks like U-Net [31] often include strong spatial compression and Transformers tend to impose high data demands for reliable generalization [32], we tailor an ultra-lightweight custom architecture that shares design elements with FireNet [33], adapted for very small frame sizes yet variable depth dimensions. Similarly to RAFT-Stereo [34], we adopt convolutions and a Gated Recurrent Unit (GRU) to efficiently handle different depth resolutions, enabling customization of the desired depth precision without modifying architecture.
@@
Finally, a dense layer that preserves the dimension of with ReLU activation and a subsequent output dense layer are applied to process the hidden state. We introduce two versions of the network for custom modification of depth estimation density. In the single-pixel version, the network predicts the normalized depth for the selected pixel while in the multi-pixel version, the output is a grid representing the normalized depth predictions of the central pixel and its 8 neighbors. Finally, normalized depth is converted into actual depth by mapping to .
@@ 3.3 Training and Inference
As supervised loss function for training the neural network model we use the mean absolute error (MAE), with given ground truth depth (this is the case of standard real-world datasets used, such as MVSEC and DSEC – see Sec.˜4). To reduce training time and further improve generalization, we additionally employ ensemble learning (EL) [36, 37]. For training, we initialize two identical but independent instances of our neural network with different random weights. The training set is split into two disjoint subsets, enabling parallel training. During testing or inference, each Sub-DSI is processed simultaneously by both networks, and the final depth estimation is obtained by averaging the individual predictions. This helps reduce variance in the predictions, leading to more stable and accurate results. We also present results without EL in Tabs.˜11, 12, 13 and 14 in the Appendix Appendix˜B.
@@ 4 Experiments
Dataset Dataset Details DSI Parameters Gauss. Filter Training Process Sequences Resolution LiDAR Span D Sub-DSI Window C Batch Optimizer LR LF Epochs MVSEC Indoor flying 346 260 px 50 ms 1 s 1 m 6.5 m 100 77 99 10 64 AdamW MAE 3 DSEC Zurich04a 640 480 px 100 ms 0.2 s 4 m 50 m 100 77 99 2 64 AdamW MAE 3
@@
We investigate the impact of DSIs derived from both monocular and stereo settings during training and testing. Table˜2 provides an overview of the key parameters used in the datasets and the training processes of the experiments. Abbreviations are: minimum depth (), maximum depth (), depth dimensions (), filter window size (Window), subtractive constant (), batch size (Batch), learning rate (LR), and loss function (LF).
@@ 4.1 Metrics
Method MVSEC DSEC Algorithm Modality Mean Err Median Err bad-pix #Points Mean Err Median Err bad-pix #Points [cm] [cm] [%] [million] [m] [m] [%] [million] SOTA EMVS [11] monocular + EMVS [11] monocular + ESVO [19] stereo SGM [38] stereo GTS [39] stereo MC-EMVS [12] stereo + MC-EMVS [12] stereo + MC-EMVS [12] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net (multi-pixel) stereo + DERD-Net (multi-pixel) stereo +
The performance of the networks is evaluated using ten standard metrics commonly employed in depth estimation tasks [12]. We calculate both mean and median errors between the estimated and GT depths, with median errors providing robustness against outliers. Additionally, we report the number of reconstructed points, reflecting the algorithm’s ability to generate valid depth estimations, and the number of outliers (bad-pix [40]), representing the proportion of significant depth estimation errors. In the Appendix, we also compute the scale-invariant logarithmic error (SILog Err) to evaluate the error while considering scale, and the sum of absolute relative differences (AErrR) to assess the relative accuracy of the depth predictions. Finally, we report -accuracy values, which indicate the percentage of points whose estimated depth falls within specified limits relative to GT [41].
@@ 4.2 Baseline Methods
We compare our approach against several SOTA methods that have been benchmarked on the task of long-term event-based depth estimation [5], thus evaluated under the same input conditions (events and camera poses) and output format (semi-dense depth maps) supportive of SLAM. In the absence of other deep stereo methods that learn from input camera poses, we also include comparisons against the SOTA instantaneous end-to-end learning-based stereo methods in Sec.˜4.5 for completeness. We adopt the same train-test splits established in prior work and standard benchmarks [5].
The Generalized Time-Based Stereovision (GTS) method [39] utilizes a two-step process: first performing stereo matching based on a time-consistency score for each event, followed by depth estimation through triangulation. The Semi-Global Matching (SGM) method [38] is adapted for event-based data by generating time images and subsequently applying stereo matching, with the depth map being refined by masking it at the locations of recent events. Another method, Event-based Stereo Visual Odometry (ESVO) [19] (ESVO2 [42]), integrates depth estimates by employing Student-t filters, ensuring robust spatio-temporal consistency between stereo time image patches.
The two closest baseline methods for performance comparison of our method are EMVS for monocular vision [11] and MC-EMVS for stereo vision [12]. Both methods extract pixel-wise depth from DSIs by applying the argmax function. To ensure consistency and fairness, we benchmark the methods following the procedure established in prior works [5].